OpenStudy (anonymous):

THE LENGTH OF A RECTANGLE IS 2 FEET LESS THAN TWICE ITS WIDTH. IF THE PERIMETER IS 26 FEET, FIND THE LENGTH AND WIDTH (PLEASE HELP) I KNOW THE ANSWER...BUT I HAVE TO SEE HOW ITS DONE.

OpenStudy (anonymous):

Perimeter = 2*l+ 2*w l=2*w-2 Perimeter = 2*(2*w-2) + 2*w =26 4w-4+2w = 26 6w = 30 w=5, l=8

OpenStudy (anonymous):

OMG that is awesome thank you!!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

I'm sorry but how did you get the 8...I see that 30/6 =5 but where did the 8 come from?

OpenStudy (anonymous):

never mind :)

OpenStudy (anonymous):

I see it

OpenStudy (anonymous):

l=2*w-2 l=2*5-2=8

OpenStudy (anonymous):

I cannot tell you how grateful i am!! i've been stuck on that for hours. :(

OpenStudy (anonymous):

Can you tell me the formula for time...when I=PRT ...how do i isolate T?

OpenStudy (anonymous):

Anytime. I don't understand. So T is time... what's P, I and R?

OpenStudy (anonymous):

So it's to figure out Simple Interest... I is Interest...P is Principle...R is Rate such as the % of interest charged and T is time... so I=P*R*T

OpenStudy (anonymous):

INTEREST IS EQUAL TO PRINCIPLE * RATE * TIME

OpenStudy (anonymous):

My problem is I need to know... For how many years must $10,000 (that's principle) be invested at an 8 1/2 % annual interest rate to yield $4,250 in simple interest?

OpenStudy (anonymous):

4250=8.5*10,000*T I think this is the equation but I'm not sure

OpenStudy (anonymous):

LOL I figured it out...heeheehehe